OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given line and has the given slope? Point= (8,-8) Slope= 3

OpenStudy (anonymous):

y-y1 = m(x-x1) Plug slope for m Plug (x,y) for (x1,y1)

OpenStudy (anonymous):

ok so the answer i got was y-8=3(x-8)

OpenStudy (anonymous):

No, plug always plug stuff in like: y - (-8) = m (x - (8))

OpenStudy (anonymous):

so is my answer wrong or do I just need parentheses?

OpenStudy (anonymous):

answer: y + 8 = 3 (x-8) it just helps you figure out the sign

OpenStudy (anonymous):

thank you!